in peas,tall (t) is dominant over dwarf (t), yellow is dominant over (y) and smooth (s) is dominant over wrinkled (s). what feraction of the offspring in the following cross would be homogygous recessive for all gene pair in the cross:yyttss × yyttsd?
1
Expert's answer
2017-10-18T12:48:07-0400
The probability for each parent to give y allele is 1/2.
Thus the probability for the first locus to be yy is:
1/2 x 1/2 = 1/4.

The probability to give allele t for the first parent is 1/2, while for the second one it is 1.
Thus, the probability for the second locus to be tt is:
1/2 x 1 = 1/2.

The probability to give allele s for the first parent is 1, while for the second one it is 1/2.
Thus, the probability for the third locus to be ss is:
1 x 1/2 = 1/2.

The probability of genotype is:
1/4 x 1/2 x 1/2 = 1/16.